A solution of 9% is mixed with a solution of 30% to form an 18% solution. How many of each solution should be used to form 21 liters of the 18% solution?
Answer by mananth(16946)

A solution of 9% is mixed with a solution of 30% to form an 18% solution. How many of each solution should be used to form 21 liters of the 18% solution?

Let 30% solution added be x
9%------------------ 30%------------------ 18% mix
21-x---------------------x-----------------------21
0.09(21-x)+0.3x= 0.18*21
1.89-0.09x+0.3x = 3.78
1.89 + 0.21x =3.78
0.21x= 3.78-1.89
x= 1.89/0.21
x= 9 liters = 30 % solution
21-9 = 12 liters 9% solution